Page tree
Skip to end of metadata
Go to start of metadata

Attendance:

D. Williams, T. Grundy, C. Mercier, J. Lesho, R. Perry.

CMS ops

Currently executing on board is STP-168, the first Short-Term Planning period of LTP05, which runs through the end of the year and into the beginning of Nominal Mission Phase.

Since the last meeting on there has been the VGAM on . See plots below.

Anomalies (if any)

Payload was switched off, just after midnight on , by Fault Detection Isolation and Recovery in response to a flooding of the SSMM by service 5 TM messages from one of the instruments (not SPICE). Recovery was attempted on 17 Sep, but the SSMM file stores' partition table was found to be corrupted, so the SSMM was cold power-cycled to return to a recoverable state. Payload recovery activities were delayed to Monday 20th September. SPICE was recovered that evening, and timeline reentered on Tuesday ~16:00 UT.

CMS data and interpretation 

See IAS CMS data tool
 
Operations times shown for CMS since the last meeting. Mainly discussed during the meeting were the measurements around and just after the VGAM.

LTP04:

  1. STP-158: 2021-07-25T22:05:00Z - 2021-07-26T07:15:00Z (DOY 206 – 207) around TCM + WOL <NORMAL>
  2. STP-159: 2021-08-01T19:05:00Z – 2021-08-02T13:25:00Z  (DOY 213 – 214) around TCM + WOL
  3. STP-159: 2021-08-04T22:30:00Z – 2021-08-05T01:30:00Z (DOY 216 – 217) around TCM + WOL
  4. STP-160: 2021-08-05T21:05:00Z – 2021-08-06T08:25:00Z (DOY 217 – 218) around TCM + WOL
  5. STP-161: 2021-08-08T19:25:00Z – 2021-08-09T08:35:00Z (DOY 220 – 221) around TCM + VGAM
    • (closest approach at 04:41).
    • thruster firings at 20:04 to 20:10, and 06:59 to 07:05
    • Decontamination heaters switched off so the temperatures would have dropped
    • Not seeing much deposition
    • On the blue trace, the HSD was closed throughout. CMS 2 temperature was dropping . CMS1 was also on a falling trend of temp  – recovery was somewhere on the 10th. Almost 24 hours of below-normal temperature.
    • Can we look at the temperature vs the frequency.
    • Thrusters look clean
    • We don't seem to have accumulated anything permanently.
    • For future, we might try to run the CMS for a day after the GAM.
    • AI: DW to ask inside SOC if we have a diagram of the thruster positions.
  6. STP-161: 2021-08-12T18:25:00Z – 23:00:00Z (DOY 224) around WOL + SA ROT 56 to 60
    1. rotation occurs 21:28:59 to 21:29:54 (within the available TM time resolution).
  7. STP-162: 2021-08-13T06:05:00Z – 21:40:00Z (DOY 225) around TCM + WOL
  8. STP-162: 2021-08-19T21:00:00Z – 00:00:00Z (DOY 231) around WOL
  9. STP-163: 2021-08-24T03:30:00Z – 09:30:00Z (DOY 236) around TCM
  10. STP-163: 2021-08-26T21:00:00Z – 00:00:00Z (DOY 238) around WOL
  11. STP-164: 2021-09-03T11:00:00Z – 14:00:00Z (DOY 246) around SA relubrication
  12. STP-165: 2021-09-06T00:05:00Z – 09:15:00Z (DOY 249) around TCM + WOL
  13. STP-165: 2021-09-09T21:00:00Z – 00:00:00Z (DOY 252) around WOL
  14. STP-166: 2021-09-15T00:05:00Z – 03:05:00Z (DOY 258) around WOL (this was the PTR test STP)
  15. STP-167: 2021-09-18T23:05:00Z – 2021-09-19T02:05:00Z (DOY 261 – 262) around WOL (see anomaly above)
  16. STP-167: 2021-09-25T23:05:00Z – 2021-09-26T02:05:00Z (DOY 268 – 269) around WOL.


Observations from the VGAM data:

  • Looks like the Thermogravimetric Analysis operations (TGAs) aren't pulling material off, so it's either stuck on and really hard to get rid of, or it's not there.
  • For the EGAM in November, it would be good to be on for 1 day before and ≤1 day after. Could lower the sample rate if needed.
  • CMS requires 2.5 W + 14 W, plus a few more watts if we actively need to keep the temperature stable. This is the time where we would want to run a fixed temperature if possible  and let it cool and heat as need be so we can ask EMC WG if that's doable.

AI: SOC to Put this to MAG & RPW.

2 campaigns have already been run with the Peltier device on. Could point RPW and MAG at the times of the TGAs that we've done in the past. If we can stay on at constant temperature, that'll make the data easier to analyse. If not, let the temperature float then the equilibration time plus the changing temperature might obliterate the data. Since we're moving into shadow, we could try holding CMS-2 warm via Peltier. Heatsink below the Mk-18 isn't going to be very low temperature, so if we heat up to 40ºC we could let it ride - we wouldn't be depositing. Could look at the data from the past and work out a good setting for this.

AI on TG, JL, SW to work out what the ops plan would be for the EGAM.


Highlights/recommendations 

    •    CMS operations to be performed weekly, capturing interesting events where possible. 
    •    CMS Operator should switch on CMS ≥30 minutes earlier than the time at which the observations would start, particularly if the observations to catch (for example) a WOL would need to start at midnight. CMS operations in SOOP Kitchen are timed such that they start 30 minutes after assumed switch-on. 
    •    Note that From LTP04 onwards (i.e., starting from STP-154), SOC has not been indicatingCMS operations in the EFECS with an RSW_EXT_SPICE. They will still be indicated with CMS “observations” in the LTP plan however. 

Future operations:

Already planned

Upcoming planned CMS Ops:

  1. STP-168: 2021-09-27T09:30:00Z – 12:30:00Z around SA relubrication 
  2. STP-169: 2021-10-04T00:05:00Z – 09:15:00Z around WOL + TCM
  3. STP-170: 2021-10-12T18:00:00Z – 21:00:00Z around SA rotation (60º → 56º)
  4. STP-170: 2021-10-13T01:00:00Z – 21:20:00Z around WOL + MAG calibration roll (2021-10-13T04:00:00Z – 18:16:40Z )
  5. STP-171: 2021-10-20T12:00:00Z – 15:00:00z around SA rotation (56º → 30º)
  6. STP-172: 2021-10-25T00:05:00Z – 03:05:00Z around WOL

Request to run CMS during the EGAM itself if power allows is still being kept alive, subject to adequate power. Unfortunately we are still awaiting power estimates from the s/c prime. SOC will push for this information in the next few days.

CMS meeting communication 

Udo had put the question about how to inform people in instrument teams about the status of the CMS measurements. As of this meeting, our minutes will be online in SOC Public at this address: CMS Ops Team meeting minutes. This is in line with those of other operational working groups (Low-latency pipeline engineering, EMC WG).

AOB

The long-term trend since March 2021 was briefly discussed. We'll update the plot here each month.


Next meeting 

Next meeting will be Thursday at 17:00 CEST / 16:00 BST / 11:00 EDT. (Daylight saving time ends in Europe on the following Sunday, and in the US on 7th Nov). This avoids clashing with the last day of the SOWG meeting for LTP06.

Advance notice, too, that for November, we will delay the meeting until after the EGAM data are back, i.e., to Thursday   17:00 CET / 16:00 GMT / 11:00 EST.




Action Items

New AIs arising from this meeting in bold.

AI-12: MS to provide CMS1 & CMS2 FoV Open (DW will ask MS) important for next VGAM
AI-13: David Williams to investigate SWA temperatures 28th Jan – Open
AI-16: OE will contact Airbus to see what inputs they need from CMS and SOC to compare with their thermal model
    Contacted ADS: Awaiting an answer when the responsible person is back from holidays.
AI-25: Rad and Orcun to take a look at the TGA data when possible, and compare with previous TGAs.
AI-27: TG, JL, SW to work out what the ops plan would be for the EGAM once the EMC constraints are known
AI-28: SOC to ask the EMC WG to look at the prior times of CMS TGAs  to see if the heater control causes problematic emissions.




  • No labels